OpenStudy (anonymous):

Find the point-slope form of the line that passes through given the point and has the given slope. (0, 0) m = 2

OpenStudy (anonymous):

@sleepyjess ..?

OpenStudy (sleepyjess):

Do you know what point slope form is?

OpenStudy (anonymous):

No. Just slope and slope-intercept form

OpenStudy (amorfide):

you want to get an equation in the for y=mx+c where m is the gradient(slope) and c is the y intercept to get the equation of the line you use an equation y-y1=m(x-x1) where y1 is the y value in the set of coordinates given, and x1 is the x value in the coordinates given substitute the values in

OpenStudy (sleepyjess):

Point-slope form: \(y-y_1 = m(x-x_1)\). \(y_1\) and \(x_1\) are the given coordinate, and m is the slope.

OpenStudy (anonymous):

@amorfide I thought the slope-intercept form was y=mx+b, not plus c.

OpenStudy (amorfide):

+c or +b same thing

OpenStudy (alexandervonhumboldt2):

it doesnt matter b it or c you can use any variabels

OpenStudy (alexandervonhumboldt2):

any letters

OpenStudy (anonymous):

Not the same thing.

OpenStudy (amorfide):

they are both constants... just different label

OpenStudy (alexandervonhumboldt2):

same thibng

OpenStudy (anonymous):

And again... Not the same thing. Is the letter "b" the exact same letter as "c"? Nope, didn't think so. Try again

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

q=w=e=r=t=y=u=i=o=p=a=s=d=f=g=h=j=k=lz=x=c=v=b=n=m

OpenStudy (alexandervonhumboldt2):

y=ax+b is also the form of slope-intr

OpenStudy (alexandervonhumboldt2):

there is no difference

OpenStudy (amorfide):

letter b is the same as letter c in the point-slope equation using the same set of coordinates and gradient I love how you are getting mad over a different representation rather than allowing help

OpenStudy (anonymous):

Not getting mad. I told you yesterday or the day before, that I don't want help from you. @amorfide